Solar Energy Corp. has $8 million in earnings with three million shares outstanding. Investment bankers think the stock can justify a P/E ratio of 25. Assume the underwriting spread is 5 percent. What should the price to the public be?
Answer #1

Price to public = P/E ratio *earnings/shares outstanding = 25*8000000/3000000= 66.67